OpenStudy (sleepyjess):

Last one ^_^ \(\sf\dfrac{sinx}{1-cosx}+\dfrac{sinx}{1+cosx}=2cscx\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \dfrac{\sin x}{1-\cos x}+\dfrac{\ sin x}{1+\cos x}\\~\\ =\sin x\left(\dfrac{1}{1-\cos x}+\dfrac{1}{1+\cos x}\right)\\~\\ =\sin x\left(\dfrac{1+\cos x+1-\cos x}{1-\cos^2 x}\right)\\~\\ =\sin x\left(\dfrac{2}{1-\cos^2 x}\right)\\~\\ =\sin x\left(\dfrac{2}{sin^2 x}\right)\\~\\ }\end{align}\)

Parth (parthkohli):

@Brostep0s mathmath333's solution is correct.

OpenStudy (mathmath333):

thanks @ParthKohli :)